ISRAEL JOURNAL OF MATHEMATICS

TABLE OF CONTENTS, VOLUME 213, 2016

Ayala, J. and H. Rubinstein The classification of homotopy classes of bounded curvature paths

79

Bader, U., B. Duchesne and J. Lécureux Amenable invariant random subgroups

399

Baik, H., S. Kim and T. Koberda Right-angled Artin groups in the C^{∞} diffeomorphism group of the real line

175

Biliotti, L., A. Ghigi and P. Heinzner Invariant convex sets in polar representations

423

Bukh, B. and L. Zhou Twins in words and long common subsequences in permutations

183

Duchesne, B. See U. Bader

Dutertre, N. Euler obstruction and Lipschitz–Killing curvatures  109

Fila, M. and M. Winkler Rate of convergence to separable solutions of the fast diffusion equation

1

Ghigi, A. See L. Biliotti Hagen, M. F. and P. Przytycki Erratum to “Cocompactly cubu-

lated graph manifolds” Israel Journal of Mathematics 207 (2015), 377–394

505

Heilman, S., E. Mossel and J. Neeman Standard simplices and pluralities are not the most noise stable

33

Heinzner, P. See L. Biliotti Jakobson, D. and F. Naud Resonances and density bounds for con-

vex co-compact congruence subgroups of SL_{2}(ℤ)

443

Kaiser, T. Global complexification of real analytic globally subanalytic functions

139

Kim, S. See H. Baik Koberda, T. See H. Baik Kwietniak, D., P. Oprocha and M. Rams On entropy of dynam-

ical systems with almost specification

475

Lécureux, J. See U. Bader Magaard, K. and G. Stroth Groups of even type which are not of

even characteristic, I

211

Magaard, K. and G. Stroth Groups of even type which are not of even characteristic, II

279

Mossel, E. See S. Heilman Naud, F. See D. Jakobson Neeman, J. See S. Heilman Oprocha, P. See D. Kwietniak Peled, R., A. Sen and O. Zeitouni Double roots of random Lit-

tlewood polynomials

55

Przytycki, P. See M. F. Hagen Rams, M. See D. Kwietniak Rubinstein, H. See J. Ayala Sen, A. See R. Peled Stroth, G. See K. Magaard

Sulca, D. Zeta functions of virtually nilpotent groups  371

Winkler, M. See M. Fila Zeitouni, O. See R. Peled Zhou, L. See B. Bukh